Here are four different comments:

(1) C (s) + H2 (g) => CH4 (g)

(2) C2H6 (g) + O2 (g) => CO2 (g) + H2O (l)

(3) C (s) + O2 (g) => CO2 (g)

(4) H2 (g) + O2 (g) => H2O (l)

A. Formulate a methane fire reaction (CH4) according to the given reactions, note that not all reactions are balanced and unnecessary Use all responses.

B. What is the theoretical amount of water you will get from the fire response you received in Section A? (In units of grams) In reaction oxygen is 124.79 grams and methane is 31.76

.C. Formulate a methane and ethane (C2H6) fire reaction according to the given reactions. What is the practical use of carbon dioxide? Oxygen? It is known that the amount of oxygen is 193.59 gram, the amount of methane is 19.25 gram and the amount of ethane is 36.47 grams and the final mole carbon dioxide is 3 mole

D. What are the possible reasons for this exploitation?

E. What are the possible reasons for a higher practical utilization than theoretical utilization?
Answer #1

A) Reaction 1 is reversed, 3 and 4 are added and the reaction is carried out:

CH4 + 2 O2 = CO2 + 2 H2O

B) The moles of the reagents are calculated:

n CH4 = g / MM = 31.74 / 16 = 2 mol

n O2 = 124.79 / 32 = 4

Both reagents are limit, the amount of water is calculated:

g H2O = 2 mol CH4 * (2 mol H2O / 1 mol CH4) * (18 g H2O / 1 mol) = 72 g

C) The reaction with ethane is:

C2H6 + CH4 + 11/2 O2 = 3 CO2 + 5 H2O

D) It can explode due to the rapid reaction speed and the release of a large amount of energy.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.